OpenStudy (anonymous):

Hannah is growing a garden in her backyard. It is in the shape of a square with side lengths of 7 feet. Her husband said that he can grow the same amount of crops in half the space. What will the side lengths of his garden measure?

OpenStudy (anonymous):

7 feet 4.95 feet 7.5 feet 14 feet

OpenStudy (harsimran_hs4):

the question is not absolutely clear...but still i will try to answer let us say the side length of her husband`s garden is x feet the area(or space ) of the garden is x^2(x times x) space of hannah`s garden is 49 ( 7 times 7) but all he needs is half the space , so x^2 = 49/2 can you now find x ?

OpenStudy (anonymous):

x²= 49/2 x²= 24.5 x= 4.95

OpenStudy (harsimran_hs4):

yes!!